What is the mole fraction HCI in an aqueous solution contains 44.2 grams of HCI and 88.6 grams of H2O?



1
Expert's answer
2022-03-04T03:13:02-0500

Molar mass HCl = 36.5 g/mol

Mol HCl in 44.2 g = 44.2 g / 36.5 g/mol = 1.211 mol

Molar mass H2O = 18 g/mol

Mol H2O in 88.6 g = 88.6 g / 18 g/mol = 4.922 mol

Total moles = 1.211 + 4.922 = 6.133 mol

Mol fraction HCl = 1.211 / 6.133 = 0.197

= 0.197